# A Multi-fidelity Estimator of the Expected Information Gain for Bayesian Optimal Experimental Design

### v1

@@ 5.1.2 Sample reuse in the inner loop
Figure 4 plots the computational cost and design-averaged correlation trade-offs under reuse, showing strengthened correlation compared to Figure 1 especially between the $u_{1}$ and $u_{2}$ and at low values of $N_{\text{in},1}$ and $N_{\text{in},2}$. Figure 5 shows the projected variance under reuse. A low-variance region appears similar to that of Figure 2, but with lower $N_{\text{in},1}$ due to the bolstered correlation between $u_{1}$ and $u_{2}$ when $N_{\text{in},1}\approx N_{\text{in},2}$, as seen in Figure 4b. The optimum is at $\text{N}_{\text{in}}=[2425,1225]$, where $N_{\text{in},1}$ has decreased and $N_{\text{in},2}$ increased from the no-reuse case, indicating that more of the computational burden has been passed from $u_{1}$ to the less costly $u_{2}$ again due to their increased correlation. Figure 6 shows the empirical estimator mean and ±$\pm 2$ standard deviation obtained from 50 repeated estimation trials using the single-fidelity NMC and optimal-$\mathrm{N}_{\text{in}}$ MF-EIG estimators under reuse. The estimator standard deviation for MF-EIG under reuse is further reduced from its no-reuse counterpart in Figure 3.
(a) Comp. cost (dashed) and correlation to $u_{0}$ (solid)
@@
Figure 5: Case 1 (reuse). Projected MF-EIG estimator variance across $\mathrm{N}_{\text{in}}$. The contour levels use a power law normalization for enhanced visualization.
Figure 6: Case 1 (reuse). Empirical estimator mean and ±$\pm 2$ standard deviation obtained from 50 repeated estimation trials using the single-fidelity NMC and optimal-$\mathrm{N}_{\text{in}}$ MF-EIG estimators.
@@ 5.1.3 Scaled noise variant
$Y_{0}$ $=g_{0}(\Theta,\xi)+g_{0}(\Theta,\xi)\circ\mathcal{E}=\left(\Theta^{3}\xi^{2}+\Theta\exp{(-|0.2-\xi|)}\right)(1+\mathcal{E}),$ (38)
$Y_{1}$ $=g_{1}(\Theta,\xi)+g_{1}(\Theta,\xi)\circ\mathcal{E}=\left(k_{1}\,\Theta^{2.5}\xi^{1.75}+\Theta\exp{(-|0.2-\xi|)}\right)(1+\mathcal{E}),$ (39)
$Y_{2}$ $=g_{2}(\Theta,\xi)+g_{2}(\Theta,\xi)\circ\mathcal{E}=\left(k_{2}\,\Theta^{2}\xi^{1.5}+\Theta\exp{(-|0.2-\xi|)}\right)(1+\mathcal{E}).$ (40)
The optimal allocation $\mathcal{A}$ is again the GISSR estimator [8]. Figure 8a shows the empirical estimator mean and ±$\pm 2$ standard deviation obtained from 50 repeated estimation trials using the single-fidelity NMC and naïve-$\text{N}_{\text{in}}$ (with reuse) estimators. Remarkable, with a seemingly minor alteration to the noise structure, the EIG trend differs substantially from the additive noise case in Figure 3, as the local maximum at $\xi=0.2$ is now completely absent. Additionally, the identical mean curves strikingly exemplifies the unbiasedness of the MF-EIG estimator. The variance reduction achieved by MF-EIG remains similar, as summarized in Figure 8b. The empirical design-averaged variance for the naïve-$\text{N}_{\text{in}}$ MF-EIG estimator is $8.04\times 10^{-5}$, compared to $1.03\times 10^{-3}$ for the single-fidelity NMC estimator, equating to a 12.86 times variance reduction.
(a) Empirical estimator mean and ±$\pm 2$ standard deviation
@@ 5.2.1 SST turbulence closure model
We focus on a two-dimensional flow over a zero-gradient flat plate, a common verification case. The computational domain is $[-0.5,2]\times[0,1]$, with inlet boundary condition on the left edge of the domain, outlet boundary conditions on the top and right edges of the domain, and adiabatic no-slip wall boundary condition at the bottom of the domain where the plate is located. The flow has Reynolds number $5\times 10^{6}$ and is modeled using steady RANS with the two-equation shear stress transport (SST) turbulence closure model [31]. This closure model blends a near-wall $k$-$\omega$ formulation with a more conventional $k$-$\epsilon$ formulation in the free-stream regime. The full closure model can be expressed using tensor notation as:
$\frac{\partial(\rho k)}{\partial t}+\frac{\partial(\rho u_{j}k)}{\partial x_{j}}$ $=P_{1}-\beta^{\ast}\rho\omega k+\frac{\partial}{\partial x_{j}}\left[(\mu+\sigma_{k}\mu_{t})\frac{\partial k}{\partial x_{j}}\right],$ (41)
$\frac{\partial(\rho\omega)}{\partial t}+\frac{(\partial\rho u_{j}\omega)}{\partial x_{j}}$ $=\frac{\rho\gamma}{\mu_{t}}P_{1}-\beta\rho\omega^{2}+\frac{\partial}{\partial x_{j}}\left[(\mu+\sigma_{\omega}\mu_{t})\frac{\partial\omega}{\partial x_{j}}\right]+2(1-F_{1})\frac{\rho\sigma_{\omega 2}}{\omega}\frac{\partial k}{\partial x_{j}}\frac{\partial\omega}{\partial x_{j}},$ (42)
@@
[…]
@@
$\mu_{t}=\frac{\rho a_{1}k}{\max(a_{1}\omega,\Omega F_{2})},$ (44)
@@
$F_{2}=\text{tanh}(\Pi^{2})\hskip 30.00005pt\text{with}\quad\Pi=\max(2\Gamma_{3},\Gamma_{1}),\quad\Gamma_{1}=\frac{500\mu}{d^{2}\omega\rho},\quad\Gamma_{3}=\frac{\sqrt{k}}{\beta^{*}\omega d},$ (45)
@@
$F_{1}=\tanh{(\Gamma^{4})}\hskip 10.00002pt\text{with}\quad$ $\Gamma=\min\left(\max(\Gamma_{1},\Gamma_{3}),\Gamma_{2}\right),$ (47)
@@
$\gamma_{1}=\frac{\beta_{1}}{\beta^{*}}-\sigma_{\omega 1}\frac{k^{2}}{\sqrt{\beta^{*}}},\quad\gamma_{2}=\frac{\beta_{2}}{\beta^{*}}-\sigma_{\omega 2}\frac{k^{2}}{\sqrt{\beta^{*}}}.$ (48)
@@
Flow solver Grid Total grid nodes Comp. cost (seconds)
@@ 5.2.2 Optimal experimental design
$Y_{m}=g_{m}(\Theta,\xi)+\mathcal{E},\quad m=1,2,3,$ (50)
@@ 5.2.3 EIG estimation
We employ three EIG estimators for this case: single-fidelity NMC, naïve-$\mathrm{N}_{\text{in}}$ MF-EIG, and optimal-$\mathrm{N}_{\text{in}}$ MF-EIG. Based on the findings from Section 5.1.2, we implement all MF-EIG estimators with inner-loop sample reuse across their utility models. We also employ the common random number technique by fixing the random seed across $\Xi$ when sampling $Z$ in the ACV outer loop. All MF-EIG estimators are then formed from a total computational budget of $w_{\text{budget}}=1.79\times 10^{7}$ seconds. Furthermore, we fix the inner-loop sample size of the high-fidelity utility model, $u_{0}$, to $N_{\text{in},0}=136$. When applying the same total budget and inner-loop sample size to the single-fidelity NMC, it allows an outer-loop sample size of $N_{\text{out}}=74$ (after rounding up).
The naïve-$\mathrm{N}_{\text{in}}$ MF-EIG estimator fixes $N_{\text{in},m}=136$ for all $m$. We employ $N_{\text{pilot}}=136$ pilot samples to estimate the sample covariance following (26) at each of the $N_{\xi}=875$ designs in $\Xi$, and then average them to obtain the design-averaged utility model covariance in (27). Since each flow solve produces $g_{m}$ for all designs simultaneously through its computational grid, the pilot phase entails only $N_{\text{pilot}}$ flow solves of each solver rather than $N_{\text{pilot}}\times N_{\xi}$. The computational cost of each utility model is estimated from the average run time over these pilot simulations. Correlation and computational cost information for the utility models are shown in Table 4. From solving (16)–(17), the optimal allocation $\mathcal{A}$ is found to be a generalized version of the MFMC estimator [33] called the ACVMF estimator [22]. The resulting total number of evaluations and total computational cost of utility models under this allocation are depicted in Figure 9. We see the allocation favors many more lower-fidelity utility model evaluations, although the total cost remains higher for the higher-fidelity models.
Utility model Underlying solver Correlation to ${u_{0}}$ Comp. cost $w_{m}$ (seconds)
@@
The optimal-$\mathrm{N}_{\text{in}}$ MF-EIG estimator optimizes $\mathrm{N}_{\text{in}}$ over an integer-valued search space of $[5,136]^{3}$ following (34). The pilot sampling process is identical to the naïve-$\mathrm{N}_{\text{in}}$ case. To accelerate the computation, we relax the optimization problem to real numbers in $[5,136]^{3}$, and solve it using the Divided Rectangles (DIRECT) [17] global optimization implementation from the SciPy [46] module. The result is then rounded to the nearest feasible integer. Correlation and computational cost information for the utility models are shown in Table 5.2.3. From solving (16)–(17), the optimal allocation $\mathcal{A}$ is found to be a generalized version of the ACVIS estimator (an even more generalized version than GISSR) called the GISMR estimator [8]. The resulting total number of evaluations and total computational cost of utility models under this allocation are depicted in Figure 10. Compared to Figure 9, even more model evaluations and computational cost are now shifted from higher- to lower-fidelity utility models, taking advantage of their correlation.
Utility
model Underlying
Correlation
Optimal
Comp. cost
$w_{m}$ (seconds)
$u_{0}$ S0 1 136 $2.42\times 10^{6}$
$u_{1}$ S1 0.993 115 $3.56\times 10^{4}$
$u_{2}$ S2 0.978 114 $3.19\times 10^{3}$
$u_{3}$ S3 0.881 29 $5.02\times 10^{1}$
Table 5: Case 2. Correlation and computational cost of utility models for the optimal-$\mathrm{N}_{\text{in}}$ MF-EIG estimator.
(a) Utility model evaluations
(b) Utility model total computational cost
Figure 10: Case 2. Total number of evaluations and total computational cost of utility models for the optimal-$\mathrm{N}_{\text{in}}$ MF-EIG estimator under its optimal allocation $\mathcal{A}$.
5.2.4 Results and discussion
Table 6 presents the MF-EIG estimator variance and variance reduction ratio relative to the single-fidelity NMC, averaged over all 875 designs across $\Xi$ and at the optimal design $\xi^{\ast}$. Projected values are predicted by MXMCPy during hyperparameter optimization, and empirical values are obtained from repeating 22 estimation trials using fresh samples. Both the naïve-$\mathrm{N}_{\text{in}}$ and optimal-$\mathrm{N}_{\text{in}}$ MF-EIG estimators improve substantially from the single-fidelity NMC, achieving more than one-order-of-magnitude (19–71 times) variance reduction. Empirical variance reduction tends to be even greater than projected by MXMCPy. Importantly, the benefit received at the optimal design, $\xi^{\ast}$, is much higher than those averaged across all designs, achieving roughly two-orders-of-magnitude (94–233 times) variance reduction. Similar to Case 1 in Section 5.1, optimal-$\mathrm{N}_{\text{in}}$ MF-EIG sees a modest 14–21% design-averaged improvement over naïve-$\mathrm{N}_{\text{in}}$, and a whopping 151% improvement at $\xi^{\ast}$.
Single-fidelity NMC Naïve-$\mathrm{N}_{\text{in}}$ Optimal-$\mathrm{N}_{\text{in}}$
Variance (projected) $3.23\times 10^{-3}$ $1.67\times 10^{-4}$ $1.40\times 10^{-4}$
Variance (empirical) $9.24\times 10^{-3}$ $1.50\times 10^{-4}$ $1.32\times 10^{-4}$
Variance at $\xi^{\ast}$ (empirical) $2.43\times 10^{-2}$ $2.58\times 10^{-4}$ $1.04\times 10^{-4}$
Variance reduction ratio (projected) 19.19 23.07
Variance reduction ratio (empirical) 61.48 70.19
Variance reduction ratio at $\xi^{\ast}$ (empirical) 94.24 233.12
Table 6: Case 2. MF-EIG variance and variance reduction ratio relative to the single-fidelity NMC estimator. Projected values are predicted by MXMCPy, and empirical values are obtained from repeating 22 estimation trials.
Both the MF-EIG estimate itself and the estimator variance reduction can vary drastically across $\Xi$. Figure 11a shows the empirical estimator mean obtained from 22 repeated estimation trials using the optimal-$\mathrm{N}_{\text{in}}$ MF-EIG estimator across a zoomed-in portion of $\Xi$ close to the flat plate; Figure 11b further plots it in semi-log $x_{1}$-axis (vertical axis) to accentuate the pattern closer to the core region of turbulent flow. Recall that the full computational domain extends to $x_{1}=1$, and these plots focus on $x_{1}\in[0,0.1]$ because regions further away all have zero EIG. The reason for zero EIG is that regions far away from the plate are not affected by, and thus do not carry information about, the turbulence phenomenon. The most informative design locations lie in the region just above the leading edge of the flat plate and near the apparent boundary layer. One possible explanation is that this region is greatly affected by the SST closure model that controls the relative balance between free-stream and near-wall turbulence formulations. The optimal design location, interestingly, resides inside the boundary layer rather than on or near the boundary layer edge.
(a) Optimal-$\mathrm{N}_{\text{in}}$ MF-EIG
(b) Optimal-$\mathrm{N}_{\text{in}}$ MF-EIG in semi-log $x_{1}$-axis
Figure 11: Case 2. Empirical estimator mean obtained from 22 repeated estimation trials using optimal-$\mathrm{N}_{\text{in}}$ MF-EIG estimator. The flow is from left to right, with an adiabatic flat plate starting at $x_{0}=0$ on the bottom of the domain triggering a turbulent boundary layer. The full computational domain extends to $x_{1}=1$, and these plots focus on $x_{1}\in[0,0.1]$ because regions further away all have zero EIG.
Figure 12 plots the variance reduction ratio across the zoomed-in portion of $\Xi$. The reduction ratio is low for regions far away from the flat plate since, as mentioned earlier, they have zero EIG, but also zero variance (i.e., all samples yield the same outcome). Consequently, these regions have no estimator variance to reduce. At the same time, reducing estimator variance for these low- or non-informative designs would not be of interest anyways. In contrast, MF-EIG performs incredibly well at the optimal design location, just off the leading edge of the flat plate but well below the apparent upper edge of the boundary layer, with the naïve-$\mathrm{N}_{\text{in}}$ and optimal-$\mathrm{N}_{\text{in}}$ MF-EIG estimators achieving 94 and 233 times variance reduction, respectively, relative to single-fidelity NMC. The MF-EIG estimators also produce considerable variance reduction around the apparent boundary layer, but the effects are less drastic.
(a) Naïve-$\mathrm{N}_{\text{in}}$ MF-EIG
(b) Optimal-$\mathrm{N}_{\text{in}}$ MF-EIG
Figure 12: Case 2. Empirical variance reduction ratio relative to single-fidelity NMC obtained from 22 repeated estimation trials using the naïve-$\mathrm{N}_{\text{in}}$ and optimal-$\mathrm{N}_{\text{in}}$ MF-EIG estimators.
The variation of MF-EIG variance reduction ratio across $\Xi$ is likely affected by two main factors. The first factor is the inherent design-variation of utility model correlations. Figure 13 shows the correlation coefficient between each $u_{m}$ and $u_{0}$ averaged over $m=1,2,3$. Very high correlation emerges in certain regions, such as just off the leading edge of the flat plate. Very low correlation takes place in other regions, such as just above the apparent edge of the boundary layer. These differences are driven by the underlying physical structures and the ability of solvers in capturing them. For example, far-field regions where flows are mostly free-stream without complex fluctuations can be easily captured by all solvers and therefore exhibit high correlation. While one generally expects high correlation to translate to high variance reduction, the high correlation in the free-stream regions do not, as seen in Figure 12. This is because their estimator variance is also zero and cannot be further reduced, as previously discussed. The second factor is the design-averaging process of the covariance matrix. Since the MF-EIG estimators are constructed based on a single covariance matrix averaged over, in this case, the entire $\Xi$, the MF-EIG hyperparameters are necessarily suboptimal in regions where the correlation deviates from the average value.
Figure 13: Case 2. Correlation coefficient between each $u_{m}$ and $u_{0}$ averaged over $m=1,2,3$.
Lastly, we consider MF-EIG in the context of discriminating designs, under the noise of MC estimation, for the purposes of finding the optimal design in (6). In this case, since the design space $\Xi$ is on the computational domain itself and each flow solve computes $g_{m}$ simultaneously across the entire $\Xi$, the EIG evaluations at multiple designs can be obtained for free if we accept to use the same $z$ samples for them—that is, the same flow solves used to estimating EIG at one design can be used for estimating EIG at any other design, without needing any new flow solves. This is the same common random number technique mentioned earlier that improves estimator correlation across $\Xi$, which aids design comparison. Even with these considerations, discriminating which design is better remains difficult for single-fidelity NMC, where the estimator variance is large relative to the difference between mean values of competing designs, as seen in Figure 5.2.3. In contrast, the variance reduction provided by MF-EIG greatly alleviates this issue, now having very little overlapping of the EIG distribution between designs. Consequently, the optimal design emerges as more of a “clear winner.”

### v2

@@ 5.1.2 Sample reuse in the inner loop
Figure 4 plots the computational cost and design-averaged correlation trade-offs under reuse, showing strengthened correlation compared to Figure 1 especially between the $u_{1}$ and $u_{2}$ and at low values of $N_{\text{in},1}$ and $N_{\text{in},2}$. Figure 5 shows the projected variance under reuse. A low-variance region appears similar to that of Figure 2, but with lower $N_{\text{in},1}$ due to the bolstered correlation between $u_{1}$ and $u_{2}$ when $N_{\text{in},1}\approx N_{\text{in},2}$, as seen in Figure 4b. The optimum is at $\text{N}_{\text{in}}=[2425,1225]$, where $N_{\text{in},1}$ has decreased and $N_{\text{in},2}$ increased from the no-reuse case, indicating that more of the computational burden has been passed from $u_{1}$ to the less costly $u_{2}$ again due to their increased correlation. Figure 6 shows the empirical estimator mean and $\pm 2$ standard deviation obtained from 50 repeated estimation trials using the single-fidelity NMC and optimal-$\mathrm{N}_{\text{in}}$ MF-EIG estimators under reuse. The estimator standard deviation for MF-EIG under reuse is further reduced from its no-reuse counterpart in Figure 3.
(a) Computational cost (dashed) and correlation to $u_{0}$ (solid)
@@
Figure 5: Case 1 (reuse). Projected MF-EIG estimator variance across different possible low-fidelity inner-loop sample sizes, $\mathrm{N}_{\text{in}}$. The contour levels use a power law normalization for enhanced visualization.
Figure 6: Case 1 (reuse). Empirical estimator mean and $\pm 2$ standard deviation obtained from 50 repeated estimation trials using the single-fidelity NMC and optimal-$\mathrm{N}_{\text{in}}$ MF-EIG estimators.
@@ 5.1.3 Scaled noise variant
(38) $Y_{0}$ $=g_{0}(\Theta,\xi)+g_{0}(\Theta,\xi)\circ\mathcal{E}=\left(\Theta^{3}\xi^{2}+\Theta\exp{(-|0.2-\xi|)}\right)(1+\mathcal{E}),$
(39) $Y_{1}$ $=g_{1}(\Theta,\xi)+g_{1}(\Theta,\xi)\circ\mathcal{E}=\left(k_{1}\,\Theta^{2.5}\xi^{1.75}+\Theta\exp{(-|0.2-\xi|)}\right)(1+\mathcal{E}),$
(40) $Y_{2}$ $=g_{2}(\Theta,\xi)+g_{2}(\Theta,\xi)\circ\mathcal{E}=\left(k_{2}\,\Theta^{2}\xi^{1.5}+\Theta\exp{(-|0.2-\xi|)}\right)(1+\mathcal{E}).$
The optimal allocation $\mathcal{A}$ is again the GISSR estimator [8]. Figure 8a shows the empirical estimator mean and $\pm 2$ standard deviation obtained from 50 repeated estimation trials using the single-fidelity NMC and naïve-$\text{N}_{\text{in}}$ (with reuse) estimators. Remarkable, with a seemingly minor alteration to the noise structure, the EIG trend differs substantially from the additive noise case in Figure 3, as the local maximum at $\xi=0.2$ is now completely absent. Additionally, the identical mean curves strikingly exemplifies the unbiasedness of the MF-EIG estimator. The variance reduction achieved by MF-EIG remains similar, as summarized in Figure 8b. The empirical design-averaged variance for the naïve-$\text{N}_{\text{in}}$ MF-EIG estimator is $8.04\times 10^{-5}$, compared to $1.03\times 10^{-3}$ for the single-fidelity NMC estimator, equating to a 12.86 times variance reduction.
(a) Empirical estimator mean and $\pm 2$ standard deviation
@@ 5.2.1 SST turbulence closure model
We focus on a two-dimensional flow over a zero-gradient flat plate, a common verification case. The computational domain is $0.5,2]\times[0,1]$, with inlet boundary condition on the left edge of the domain, outlet boundary conditions on the top and right edges of the domain, and adiabatic no-slip wall boundary condition at the bottom of the domain where the plate is located. The flow has Reynolds number $5\times 10^{6}$ and is modeled using steady RANS with the two-equation shear stress transport (SST) turbulence closure model [31]. This closure model blends a near-wall $k$-$\omega$ formulation with a more conventional $k$-$\epsilon$ formulation in the free-stream regime. The full closure model can be expressed using tensor notation as:
(41) $\frac{\partial(\rho k)}{\partial t}+\frac{\partial(\rho u_{j}k)}{\partial x_{j}}$ $=P_{1}-\beta^{\ast}\rho\omega k+\frac{\partial}{\partial x_{j}}\left[(\mu+\sigma_{k}\mu_{t})\frac{\partial k}{\partial x_{j}}\right],$
(42) $\frac{\partial(\rho\omega)}{\partial t}+\frac{(\partial\rho u_{j}\omega)}{\partial x_{j}}$ $=\frac{\rho\gamma}{\mu_{t}}P_{1}-\beta\rho\omega^{2}+\frac{\partial}{\partial x_{j}}\left[(\mu+\sigma_{\omega}\mu_{t})\frac{\partial\omega}{\partial x_{j}}\right]+2(1-F_{1})\frac{\rho\sigma_{\omega 2}}{\omega}\frac{\partial k}{\partial x_{j}}\frac{\partial\omega}{\partial x_{j}},$
@@
(43) $P_{1}=\tau_{ij}\frac{\partial u_{i}}{\partial x_{j}}\hskip 30.00005pt$ $\text{with}\quad\tau_{ij}=\mu_{t}\left(2S_{ij}-\frac{2}{3}\frac{\partial u_{k}}{\partial x_{k}}\delta_{ij}\right)-\frac{2}{3}\rho k\delta_{ij}$
$\text{and}\quad S_{ij}=\frac{1}{2}\left(\frac{\partial u_{i}}{\partial x_{j}}+\frac{\partial u_{j}}{\partial x_{i}}\right),$
@@
(44) $\mu_{t}=\frac{\rho a_{1}k}{\max(a_{1}\omega,\Omega F_{2})},$
@@
(45) $F_{2}=\text{tanh}(\Pi^{2})\hskip 30.00005pt\text{with}\quad\Pi=\max(2\Gamma_{3},\Gamma_{1}),\quad\Gamma_{1}=\frac{500\mu}{d^{2}\omega\rho},\quad\Gamma_{3}=\frac{\sqrt{k}}{\beta^{*}\omega d},$
@@
(47) $F_{1}=\tanh{(\Gamma^{4})}\hskip 30.00005pt\text{with}\quad$ $\Gamma=\min\left(\max(\Gamma_{1},\Gamma_{3}),\Gamma_{2}\right),$
@@
(48) $\gamma_{1}=\frac{\beta_{1}}{\beta^{*}}-\sigma_{\omega 1}\frac{k^{2}}{\sqrt{\beta^{*}}},\quad\quad\gamma_{2}=\frac{\beta_{2}}{\beta^{*}}-\sigma_{\omega 2}\frac{k^{2}}{\sqrt{\beta^{*}}}.$
@@
Flow solver Grid Total grid nodes Computational cost (seconds)
@@ 5.2.2 Optimal experimental design
(50) $Y_{m}=g_{m}(\Theta,\xi)+\mathcal{E},\quad m=1,2,3,$
@@ 5.2.3 EIG estimation
We employ three EIG estimators for this case: single-fidelity NMC, naïve-$\mathrm{N}_{\text{in}}$ MF-EIG, and optimal-$\mathrm{N}_{\text{in}}$ MF-EIG. Based on the findings from Section 5.1.2, we implement all MF-EIG estimators with inner-loop sample reuse across their utility models. We also employ the common random number technique by fixing the random seed across $\Xi$ when sampling $Z$ in the ACV outer loop. All MF-EIG estimators are then formed from a total computational budget of $w_{\text{budget}}=1.79\times 10^{7}$ seconds. Furthermore, we fix the inner-loop sample size of the high-fidelity utility model, $u_{0}$, to $N_{\text{in},0}=136$. This choice balances bias control with our total computational budget, while enabling a reasonable outer-loop sample size if using standard NMC. Under the same budget and inner-loop configuration, the single-fidelity NMC baseline allows for an outer-loop sample size of $N_{\text{out}}=74$ (after rounding up).
The naïve-$\mathrm{N}_{\text{in}}$ MF-EIG estimator fixes $N_{\text{in},m}=136$ for all $m$. We employ $N_{\text{pilot}}=136$ pilot samples to estimate the sample covariance following (26) at each of the $N_{\xi}=875$ designs in $\Xi$, and then average them to obtain the design-averaged utility model covariance in (27). Since each flow solve produces $g_{m}$ for all designs simultaneously through its computational grid, the pilot phase entails only $N_{\text{pilot}}$ flow solves of each solver rather than $N_{\text{pilot}}\times N_{\xi}$. The computational cost of each utility model is estimated from the average run time over these pilot simulations. Correlation and computational cost information for the utility models are shown in Table 5.2.3. From solving (16)–(17), the optimal allocation $\mathcal{A}$ is found to be a generalized version of the MFMC estimator [33] called the ACVMF estimator [22]. The MXMCPy tool takes 10.89 seconds to find the optimal MF-EIG hyperparameters using a single 3.0 GHz Intel Xeon Gold 6154 processor, a negligible cost in comparison to the costs of utility model evaluations. The resulting total number of evaluations and total computational cost of utility models under this allocation are depicted in Figure 9. We see the allocation favors many more lower-fidelity utility model evaluations, although the total cost remains higher for the higher-fidelity models.
model
$w_{m}$ (seconds)
@@
The optimal-$\mathrm{N}_{\text{in}}$ MF-EIG estimator optimizes $\mathrm{N}_{\text{in}}$ over an integer-valued search space of $5,136]^{3}$ following (34). The pilot sampling process is identical to the naïve-$\mathrm{N}_{\text{in}}$ case. To accelerate the computation, we relax the optimization problem to real numbers in $5,136]^{3}$, and solve it using the Divided Rectangles (DIRECT) [17] global optimization implementation from the SciPy [46] module. The result is then rounded to the nearest feasible integer. Correlation and computational cost information for the utility models are shown in Table LABEL:table:logratio_models_opt. From solving (16)–(17), the optimal allocation $\mathcal{A}$ is found to be a generalized version of the ACVIS estimator (an even more generalized version than GISSR) called the GISMR estimator [8]. The resulting total number of evaluations and total computational cost of utility models under this allocation are depicted in Figure LABEL:fig:opt_cost_evals. Compared to Figure 9, even more model evaluations and computational cost are now shifted from higher- to lower-fidelity utility models, taking advantage of their correlation.